### v1

# Photometry and Spectroscopy of SN 2024pxl: A Luminosity Link Among Type Iax Supernovae

Abstract
We present extensive ultraviolet to optical photometric and optical to near-infrared (NIR) spectroscopic follow-up observations of the nearby intermediate-luminosity (16.810.19 mag) Type Iax supernova (SN) 2024pxl in NGC 6384. SN 2024pxl exhibits a faster light curve evolution than the high-luminosity members of this class, and slower than low-luminosity events. The observationally well-constrained rise time of 10 days and an estimated synthesized 56Ni mass of 0.03 M⊙, based on analytical modeling of the pseudobolometric light curve, are consistent with models of the weak deflagration of a carbon-oxygen white dwarf. Our optical spectral sequence of SN 2024pxl shows weak Si II lines and spectral evolution similar to other high-luminosity Type Iax SNe, but also prominent early-time C II line, like lower-luminosity Type Iax SNe. The late-time optical spectrum of SN 2024pxl closely matches that of SN 2014dt, and its NIR spectral evolution aligns with those of other well-studied, high-luminosity Type Iax SNe. The spectral-line expansion velocities of SN 2024pxl are at the lower end of the Type Iax SN velocity distribution, and the velocity distribution of iron-group elements compared to intermediate-mass elements suggests that the ejecta are mixed on large scales, as expected in pure deflagration models. SN 2024pxl exhibits characteristics intermediate between those of high-luminosity and low-luminosity Type Iax SNe, further establishing a link across this diverse class.
1 Introduction
Type Iax supernovae (SNe) are the fainter, peculiar cousins of Type Ia SNe, both arising from the explosive thermonuclear fusion of carbon-oxygen (CO) white dwarfs (WDs; for a review, see Jha, 2017). Owing to their homogeneity, “normal” Type Ia SNe serve as standardizable candles in cosmology (Phillips, 1993; Phillips et al., 1999). Type Iax SNe, however, display comparatively lower luminosities, explosion energies, and ejecta velocities (Li et al., 2003; Filippenko, 2003; Foley et al., 2013; Jha, 2017). Their physical properties are also more heterogeneous, with luminosities ranging from mag (Karambelkar et al., 2021) to mag (Stritzinger et al., 2015), and line absorption velocities at maximum light between 2000 and 8000 km s-1 (Foley et al., 2009; Stritzinger et al., 2014). These distinct characteristics suggest that they arise from progenitor systems and/or have explosion mechanisms that differ in detail from normal Type Ia SNe.
The light curves of Type Iax SNe differ markedly from those of normal Type Ia SNe, rising faster and declining more rapidly in bluer bands (Magee et al., 2016, 2017; Jha, 2017; Li et al., 2018). The light-curve-shape variations and the observed diversity in brightness (McClelland et al., 2010; Narayan et al., 2011; Foley et al., 2013; Magee et al., 2016; Singh et al., 2023) suggest that the synthesized 56Ni, powering the light curve, spans a wide range from 810-4 M⊙ (Karambelkar et al., 2021) to 0.3 M⊙ (Stritzinger et al., 2015).
Spectroscopically, Type Iax SNe resemble 91T-like Type Ia SNe (Phillips et al., 2007) at early times, characterized by strong lines due to Fe III and Fe II and typically weak Si II. As they evolve, their late-time spectra diverge significantly from those of Type Ia SNe, displaying both permitted and forbidden lines of Fe and Ca (McCully et al., 2014a; Stritzinger et al., 2015). Notably, no fully nebular spectrum of a Type Iax SN has been observed, owing to the persistence of lines with a pronounced P-Cygni profile for hundreds of days (e.g., Camacho-Neves et al., 2023). Their spectral evolution slows considerably after 200400 days past maximum light (Foley et al., 2016). Maeda & Kawabata (2022) discussed the existence of an Fe-rich innermost region and its association with a bound WD (remnant) using the day 500 spectrum of SN 2019muj.
High-luminosity and low-luminosity Type Iax SNe differ spectroscopically. For example, high-luminosity objects do not show a strong C II feature at 6580 Å, whereas low-luminosity ( mag) Type Iax SNe do. Low-luminosity objects such as SNe 2008ha (Foley et al., 2009), 2010ae (Stritzinger et al., 2014), 2019gsc (Srivastav et al., 2020; Tomasella et al., 2020), 2020kyg (Srivastav et al., 2022; Singh et al., 2023), and 2021fcg (Karambelkar et al., 2021) also exhibit rapid spectroscopic evolution, lower expansion velocities, and a faster transition to the partial nebular phase.
Weak deflagrations of CO WDs are promising explosion models for Type Iax SNe, roughly reproducing many observed properties of brighter objects, such as their 56Ni masses, peak luminosities, rise times, and early-time spectra (Fink et al., 2014; Kromer et al., 2013; Branch et al., 2004; Jordan et al., 2012; Lach et al., 2022). However, these models struggle to match low-luminosity Type Iax SNe that may instead be explained by weak deflagrations in hybrid carbon-oxygen-neon (CONe) WDs (Meng & Podsiadlowski, 2014; Kromer et al., 2015).
SN 2012Z offered a valuable opportunity to investigate the progenitor systems of Type Iax SNe. Using deep Hubble Space Telescope (HST) pre-explosion images of its host galaxy NGC 1309, McCully et al. (2014b) detected a luminous blue source at the location of the SN and proposed a progenitor system consisting of a CO WD with a helium-star companion. Follow-up observations of SN 2012Z confirmed that the SN is still brighter than pre-explosion, suggesting that the helium-star companion survived the explosion and that a bound remnant was left behind (McCully et al., 2022; Schwab et al., 2025). A similar progenitor scenario was suggested for SN 2014dt (Foley et al., 2015). Moreover, a faint red source was identified in HST images taken 4 yr after the explosion of SN 2008ha that may be a companion star or remnant (Foley et al., 2014). In the case of SN 2020udy, another bright Type Iax, constraints on interaction with a companion star from light curves obtained soon after the explosion also favor a helium-star companion (Maguire et al., 2023).
SN 2024pxl (, , J2000) was first discovered and reported to the Transient Name Server111https://www.wis-tns.org by the BTSbot machine-learning model (Rehemtulla et al., 2024b) on 23 July 2024 at 09:41:18 (UTC dates are used throughout this paper) using data from the Zwicky Transient Facility (ZTF; Bellm et al., 2019a, b; Dekany et al., 2020; Masci et al., 2019; Graham et al., 2019). As part of the BTSbot-nearby program (Rehemtulla et al., 2025), BTSbot immediately and autonomously triggered a target-of-opportunity photometric and spectroscopic request to the spectral energy distribution machine spectrograph (SEDM; Blagorodnova et al., 2018; Kim et al., 2022). SN 2024pxl is located 455 E and 78 N from the nucleus of its host galaxy, NGC 6384. Other designations of this SN are ATLAS24lpk and ZTF24aawrofs. Smartt et al. (2024) classified SN 2024pxl as a Type Iax SN, noting a strong resemblance to the bright Type Iax SN 2005hk one week before maximum light.
This study presents an extensive follow-up campaign of SN 2024pxl, integrating observations from both ground- and space-based telescopes. SN 2024pxl is perhaps the best-sampled Type Iax SN to date in terms of its photometric and spectroscopic data. Kwok et al. (2025) analyze optical and near-infrared (NIR) spectra from this study together with mid-infrared (MIR) data from the James Webb Space Telescope (JWST), providing key insights into the explosion mechanism and progenitor system, complementing the results presented here. Section 2 details the observations and data reduction, and Section 3 describes the distance determination, line-of-sight extinction, and explosion epoch estimation. Section 4 presents the light curve, color evolution, and analytical modeling of the pseudobolometric light curve. In Section 5, we examine the spectral features and line-velocity evolution of SN 2024pxl and compare it with other well-studied Type Iax SNe. Section 6 summarizes the key findings of this study.
2 Observations and Data Reduction
2.1 Photometry
A high-cadence dataset for SN 2024pxl was acquired with both ground- and space-based telescopes as part of an extensive photometric observing campaign initiated on the day of discovery. Photometry in bands was obtained through the Global Supernova Project (GSP) collaboration using the Las Cumbres Observatory (LCO; Brown et al. 2013) 0.4 m and 1 m telescopes. Preprocessing, including bias correction and flat fielding, was handled by the BANZAI pipeline (McCully et al., 2018). Further data reduction was carried out using lcogtsnpipe (Valenti et al., 2016a), a photometric reduction pipeline that uses point-spread-function (PSF) photometry (Stetson, 1987) to calculate zero points, color terms, and extracted magnitudes. Photometry in bands is reported in Vega magnitudes (Stetson, 2000), while band data are presented in AB magnitudes (Oke & Gunn, 1983), calibrated against Sloan Digital Sky Survey (SDSS) sources (Smith et al., 2002). Given the proximity of SN 2024pxl to its host galaxy, host contamination was mitigated by subtracting template images acquired on March 7, 2018 — before the explosion. These templates, originally taken for SN 2017drh (which exploded in the same galaxy), were subtracted using the PyZOGY algorithm (Zackay et al., 2016; Guevel & Hosseinzadeh, 2017), integrated within the lcogtsnpipe pipeline.
Additional imaging of SN 2024pxl was obtained in the bands with the 1 m Nickel telescope at Lick Observatory. The images were calibrated using bias and sky flat-field frames following standard procedures. PSF photometry was performed and calibrated relative to Pan-STARRS photometric standards (Flewelling et al., 2020). Optical photometry of SN 2024pxl is presented in LABEL:tab:photometric_observational_log_2024pxl.
SN 2024pxl was also observed in the z band with the RC80 and BRC80 robotic telescopes at Piszkesteto station of Konkoly Observatory and at Baja Observatory of University of Szeged, Hungary (a complete BgVriz light curve will be presented by Banhidi et al., in prep.). The magnitudes were computed via aperture photometry and tied to Pan-STARRS photometry of local stars, including a color term in the transformation. The z-band photometry of SN 2024pxl is provided in LABEL:tab:z_band_photometry_24pxl.
The Neil Gehrels Swift Observatory (Swift Gehrels et al., 2004) observed SN 2024pxl with its Ultra-Violet/Optical Telescope (UVOT; Roming et al. 2005) in both ultraviolet (UV) and optical filters. We reduced the UVOT images using the High-Energy Astrophysics Software (HEASoft222https://heasarc.gsfc.nasa.gov/docs/software/heasoft/). A circular source region centered at the position of the SN with a radius of 3 was used for aperture photometry. We measured the background contribution from a circular region (aperture radius of 5) that is not contaminated by any other sources. Zeropoints for photometry were chosen from Breeveld et al. (2010) with time-dependent sensitivity corrections updated in 2020. Table 5 details the photometric observations of SN 2024pxl collected by Swift.
2.2 Spectroscopy
2.2.1 Optical Spectra
We have compiled an extensive spectral dataset for SN 2024pxl using multiple ground-based telescopes in the optical and near-infrared (NIR) domains. Spectra of SN 2024pxl obtained with the Double Beam Spectrograph (DBSP; Oke & Gunn 1982) mounted on the 5 m Hale Telescope at Palomar Observatory were reduced using DBSP-DRP (Mandigo-Stoba et al., 2021, 2022a, 2022b) and PypeIt (Prochaska et al., 2020). Spectra acquired with the Wide Field Spectrograph (WiFeS; Dopita et al. 2007, 2010) on Australian National University (ANU) 2.3 m telescope at Siding Spring Observatory were reduced with the data-reduction pipeline PyWiFeS (Childress et al., 2014; Carr et al., 2024). Spectroscopic observations were also triggered with the FLOYDS spectrograph on the 2 m Faulkes Telescope North and South (FTN and FTS; Brown et al. 2013) through the GSP collaboration. The spectra were reduced with the floydsspec333https://www.authorea.com/users/598/articles/6566 pipeline, using standard reduction techniques.
Gemini IRAF444IRAF is distributed by the National Optical Astronomy Observatories, operated by the Association of Universities for Research in Astronomy, Inc., under a cooperative agreement with the National Science Foundation. packages were used to reduce the spectra of SN 2024pxl from the Gemini Multi-Object Spectrograph mounted at Gemini North (GMOS-N; Hook et al. 2004) and Gemini South (GMOS-S) Observatories. Two spectra of SN 2024pxl were acquired using the SPectrograph for the Rapid Acquisition of Transients (SPRAT; Piascik et al. 2014), mounted on the Liverpool Telescope (LT; Steele & et al. 2004). Reduction of the SPRAT spectra was carried out using the SPRAT pipeline.
SN 2024pxl was also observed with the Robert Stobie Spectrograph (RSS) attached to the 9.2 m Southern African Large Telescope (SALT). All RSS spectra were reduced using a custom pipeline based on standard Pyraf (Science Software Branch at STScI, 2012) routines and the PySALT package (Crawford et al., 2010). A few spectra of SN 2024pxl were taken from the Marcario Low-Resolution Spectrograph (LRS) mounted on Hobby–Eberly Telescope (HET) located in McDonald Observatory (Ramsey et al., 1998). These were reduced following standard IRAF packages.
As part of our observing campaign, we also utilized the Hanle Faint Object Spectrograph and Camera (HFOSC) mounted on the Himalayan Chandra Telescope (HCT; Prabhu & Anupama 2010). The HCT spectra were reduced using standard IRAF routines.
SN 2024pxl was also observed with Binospec on the MMT telescope (Fabricant et al., 2019); these spectra were reduced using the Binospec IDL pipeline (Kansky et al., 2019). A spectrum was taken with the Low Resolution Imaging Spectrometer (LRIS; Oke et al., 1995) on the Keck I 10 m telescope, which was reduced in a standard way using the LPipe pipeline (Perley, 2019). Many spectra of SN 2024pxl were obtained with the Kast double spectrograph (Miller & Stone, 1993) on the Shane 3 m telescope at Lick Observatory, either at low airmass or with the slit aligned along the parallactic angle to minimize the effects of atmospheric dispersion (Filippenko, 1982). The reduction of all Kast spectra was performed following standard procedures555https://github.com/msiebert1/UCSC_spectral_pipeline,666https://github.com/ishivvers/TheKastShiv as outlined by Siebert et al. (2019) and Silverman et al. (2012).
In our observing campaign, we also acquired data from the Gran Telescopio Canarias (GTC) with OSIRIS and EMIR, and reduced the spectra using a dedicated pipeline based on PypeIt (Prochaska et al., 2020a, b, 2021; Galbany et al., 2025). PypeIt was also used to reduce the spectra of SN 2024pxl gathered with the 2.56 m Nordic Optical Telescope (NOT) deploying the Alhambra Faint Object Spectrograph and Camera (ALFOSC). We collected one spectrum of SN 2024pxl with the XShooter spectrograph mounted on the Very Large Telescope (Vernet et al., 2011); the EsoReflLex pipeline (Freudling et al., 2013), including the XShooter module, was used to reduce it.
The reduced spectra were scaled to the photometry of the corresponding epoch using a linear fit. All spectra were dereddened and corrected for host-galaxy redshift. A log of the spectroscopic observations of SN 2024pxl is given in LABEL:tab:spectroscopic_observations_24pxl.
2.2.2 NIR Spectra
The first NIR spectrum in our observing campaign of SN 2024pxl was obtained from Near-InfraRed Echellette Spectrometer (NIRES) on the Keck II 10 m telescope through the Keck Infrared Transient Survey (KITS) program and reduced using procedures outlined by Tinyanont et al. (2024). Two spectra were obtained by the Folded-port InfraRed Echellette Spectrograph (FIRE; Simcoe et al., 2013) mounted on the 6.5 m Magellan-Baade Telescope at Las Campanas Observatory in Chile, and reduced using the IDL pipeline firehose (Simcoe et al., 2013). NIR spectra of SN 2024pxl were acquired from the NASA InfraRed Facility Telescope (IRTF) using the SpeX spectrograph (Rayner et al., 2003). Data reduction of SpeX data was performed with Spextool and included tasks such as flat fielding, wavelength calibration, background subtraction, and spectra extraction (Cushing et al., 2004).
We also obtained two NIR spectra of SN 2024pxl with the Espectrógrafo Multiobjeto Infra-Rojo spectrograph (EMIR; Garzón et al. 2022) mounted on the GTC, which were reduced using a dedicated pipeline based on PyEMIR (Pascual et al., 2010; Cardiel et al., 2019). The Southern Astrophysical Research (SOAR) telescope, equipped with TripleSpec (Schlawin et al., 2014), was also deployed to obtain NIR spectra. Spectra acquired by SOAR were reduced using the IDL-based Spextool package (Cushing et al., 2004). All NIR observations for SN 2024pxl are reported in Table 7.
3 Distance, extinction, and explosion epoch
SN 2024pxl is hosted by NGC 6384 at a spectroscopic redshift of (Lagattuta et al., 2013). Fortunately, a normal (though highly dust-extinguished) SN Ia with a well-sampled light curve from LCO and Lick Observatory, SN 2017drh, was also hosted by NGC 6384 (Stahl et al., 2019). We estimate the distance to SN 2017drh using BayeSN, a hierarchical Bayesian SN Ia light-curve model (Mandel et al., 2022; Grayling et al., 2024). SN 2017drh was located near the center of NGC 6384 and is significantly dust reddened, much more than SNe Ia typically used for cosmological analyses. Thus, we favored using BayeSN because it explicitly models host-galaxy dust reddening and extinction separately from intrinsic SN spectral energy distribution (SED) variations, unlike alternatives like SALT2 (Guy et al., 2007) that aggregate intrinsic and extrinsic SN Ia color variations.
Using the BayeSN model trained by Ward et al. (2023), and assuming an reddening law, fitting the light curve of SN 2017drh yields a distance modulus to NGC 6384 of mag, which corresponds to a distance of Mpc that we adopt for our analysis of SN 2024pxl. The BayeSN fit to SN 2017drh results in a host-galaxy extinction of mag, confirming the large amount of dust along the line of sight to SN 2017drh. We present the LCO photometry, BayeSN fit, and spectroscopic observations of SN 2017drh in Appendix A.
The extinction experienced by SN 2024pxl due to the Milky Way in the direction of NGC 6384 is mag, corresponding to mag (Schlafly & Finkbeiner, 2011) and assuming = 3.1. Estimating the host-galaxy extinction is more difficult because, unlike the normal SN Ia 2017drh, we do not have a standard model color for the diverse Type Iax class. However, in our spectroscopic observations of SN 2024pxl with sufficient spectral resolution, we note the presence of narrow interstellar Na I D absorption lines from both the Milky Way and NGC 6384, with approximately equal strength. Thus, we crudely infer the host-galaxy contribution to the extinction of SN 2024pxl to be similar to the Milky Way extinction and adopt a total reddening and extinction along the line of sight to SN 2024pxl of mag and mag.
A nondetection of SN 2024pxl was reported on 21 July 2024 (JD = 2460512.75; Rehemtulla et al., 2024a) with a limiting AB magnitude of 19.37 (3 upper limit) in the r filter. To estimate the explosion epoch of SN 2024pxl, we combined photometric data from the ATLAS and ZTF surveys in four optical bands (ATLAS c and o; ZTF g and r). A power-law model of the form
was used to fit the early-time flux evolution, where is a band-dependent scaling constant, is the explosion time, and is the power-law index.
The model was simultaneously fit to all the observed fluxes in different bands (Figure 1). Unlike the commonly adopted fixed assumption used in early light-curve fits to model a homologously expanding fireball (e.g., Arnett 1982; Nugent et al. 2011), we allowed to vary as a free parameter to account for deviations observed in some thermonuclear SN subclasses (e.g., Magee et al. 2016). The best-fit result yields a common explosion epoch of
with a power-law index of
indicating a shallower rise than the canonical case.
4 Light-Curve Properties
4.1 Light Curves and Color Curves
Figure 2 presents the evolution of the light curve for SN 2024pxl in the BgVriz bands, with dense sampling around maximum brightness in all bands. We use a low-order polynomial fit to estimate the time and magnitude at maximum light in the BgVriz bands. Additionally, we estimate the decline in magnitude from the light-curve peak to 15 days after () for these bands. These light-curve parameters for SN 2024pxl are listed in Table 1. Figure 3 displays the evolution of SN 2024pxl in the Swift UVOT bands. All bands shown in Figure 3, except for UVW2 and UVM2, cover the peak well. We estimate the Swift UVOT light-curve data parameters in Table 2.
| SN 2024pxl | band | band | band | band | band | band |
| JD of maximum light (2,460,000+) | 524.60.5 | 525.20.5 | 527.40.5 | 528.140.5 | 530.50.5 | 531.70.5 |
| Magnitude at maximum (mag) | 16.420.02 | 16.120.01 | 15.910.02 | 15.580.01 | 15.500.01 | 15.470.07 |
| Absolute magnitude at maximum (mag) | 0.19 | 0.19 | 0.19 | 0.19 | 0.19 | 0.21 |
| (mag) | 1.790.09 | 1.510.04 | 0.840.06 | 0.480.04 | 0.530.04 | 0.490.10 |
| SN 2024pxl | UVW2 | UVM2 | UVW1 | Swift | Swift | Swift |
| JD of maximum light (2,460,000+) | – | – | 521.50.5 | 521.90.5 | 525.30.5 | 527.80.5 |
| Magnitude at maximum (mag) | – | – | 17.830.14 | 16.210.08 | 16.460.06 | 15.900.07 |
| Absolute magnitude at maximum (mag) | – | – | 0.24 | 0.20 | 0.20 | 0.20 |
| (mag) | – | – | 1.460.61 | 2.170.32 | 1.780.15 | 0.950.12 |
To further explore the nature of the light-curve evolution of SN 2024pxl, we compare its BgVri light curves with those of several well-studied Type Iax SNe (Figure 4). For this, we choose high-luminosity Type Iax SNe 2002cx (Li et al., 2003), 2005hk (Phillips et al., 2007; Sahu et al., 2008), 2011ay (Szalai et al., 2015), 2012Z (Stritzinger et al., 2015), 2020rea (Singh et al., 2022), and 2020udy (Singh et al., 2024); intermediate-luminosity Type Iax SN 2019muj (Barna et al., 2021); and low-luminosity Type Iax SNe 2008ha (Foley et al., 2009), 2010ae (Stritzinger et al., 2014), and 2019gsc (Srivastav et al., 2020; Tomasella et al., 2020). These comparison SNe cover both ends of the luminosity distribution for Type Iax SNe. The rest-frame magnitudes for each SN displayed in Figure 4 are normalized to the peak in the respective bands.
In the optical bands, SN 2024pxl consistently exhibits decline rates that are faster than those of the high-luminosity Type Iax SNe and slower than those of the low-luminosity events, placing it between these two groups in terms of photometric evolution. Its behavior is broadly similar to that of the intermediate-luminosity SN 2019muj, although SN 2024pxl evolves somewhat more slowly in several filters. This suggests that SN 2024pxl lies between SN 2019muj and high-luminosity Type Iax SNe in optical decline rates.
In the UV, SN2024pxl displays a slow decline relative to other Type Iax SNe, with a rate comparable to that of a normal Type Ia SN (Milne et al., 2010). This distinguishes its UV behavior from the luminosity-dependent trend observed in the optical bands. A comparison with Swift UVW1 light curves of SNe 2005hk, 2011ay, 2012Z, 2019muj, 2020kyg, and 2020udy is presented in Figure 5.
Figure 6 illustrates the evolution of SN 2024pxl’s colors (B-V, V-I, V-R, and R-I) in comparison with other Type Iax SNe. All colors have been adjusted for total reddening. The color evolution of SN 2024pxl closely resembles that observed in other Type Iax SNe selected for comparison.
4.2 Bolometric Light Curve
The pseudobolometric BgVriz light curve of SN 2024pxl is generated using SuperBol (Nicholl, 2018), incorporating the distance and extinction values outlined in Section 3. SuperBol takes the dereddened fluxes and uncertainties across passbands to construct an SED at each epoch. Luminosities are determined by integrating this spectrum while propagating uncertainties. To account for missing passbands in the UV and NIR, SuperBol fits a blackbody to each SED to derive a correction which is then applied to the pseudobolometric luminosities.
The pseudobolometric light curves of the other comparison SNe presented in Figure 7 are constructed using a similar approach. SN 2018cni (Singh et al., 2023), another bright Type Iax, is also included for comparison in Figure 7. The peak pseudobolometric BgVriz luminosity of SN 2024pxl is erg s-1. SN 2024pxl is less luminous than all the high-luminosity Type Iax SNe, such as SNe 2012Z, 2020rea, and 2020udy, but is brighter than the intermediate-luminosity Type Iax SN 2019muj. SN 2024pxl lies in the gap between the high-luminosity and low-luminosity Type Iax SNe, in agreement with our decline-rate comparison (Section 4.1).
We conduct analytical modeling of the pseudobolometric light curve for SN 2024pxl using the methods outlined by Arnett (1982) and Valenti et al. (2008). The basic assumptions of this model consist of a small initial radius, constant optical opacity, spherically symmetric and optically thick ejecta, and the inclusion of 56Ni in the ejected material. By fitting the pseudobolometric light curve of SN 2024pxl, we estimate a 56Ni mass of M⊙, an ejecta mass of M⊙, and a rise time of 10 days, with uncertainties derived from the covariance matrix. For the fit, we assume a constant opacity of cm2 g-1 and a photospheric velocity of 5000 km s-1 at maximum brightness. Directly estimating the photospheric expansion velocity from the bottom panel of Figure 8, we calculate km s-1 over the nearly linear expansion period from maximum light to 15 days. It is surprising that this value is larger than the typical photospheric absorption-line velocities, perhaps suggesting that the optical depth approaches unity farther out in the ejecta than the lines would predict.
We also compared the evolution of the blackbody temperature and radius of SN 2024pxl with those of other Type Iax SNe (Figure 8). The blackbody temperature of SN 2024pxl is lower than that of the different comparison SNe before maximum light. After that, it remains at the lower end of the blackbody temperature distribution. It is possible that this could be the effect of underestimated extinction. The evolution of the blackbody radius of SN 2024pxl is proportional to the luminosity of the SN, which is positioned in an intermediate location in Figure 8. SN 2024pxl’s blackbody temperature and radius evolution align well with the characteristic patterns seen in other Type Iax SNe.
For Type Iax SNe, contributions from the unobserved UV and IR passbands remain poorly constrained. However, based on available data for a few Type Iax SNe, the combined UV and NIR contributions to the total bolometric flux have been estimated. Phillips et al. (2007) demonstrated that the UV flux accounts for approximately 20% of the UV-through-IR (UVOIR) light curve for an early photospheric phase in the case of SN 2005hk. Similarly, Yamanaka et al. (2015) estimated that the NIR band makes up about 20% of the optical+NIR flux around peak light for SN 2012Z. Tomasella et al. (2016) found that the UV+IR flux collectively accounted for % of the total emission in SN 2014ck. Moreover, Srivastav et al. (2020) and Dutta et al. (2022) calculated the ratio of (peak pseudobolometric luminosity) to (peak blackbody luminosity) as 0.69 for SN 2019gsc and 0.62 for SN 2020sck, respectively. In the case of SN 2020kyg, Srivastav et al. (2022) reported that about 60% of the total bolometric luminosity near maximum light is attributed to the optical emission.
We have Swift observations that capture the light-curve evolution of SN 2024pxl around the peak. Using SuperBol for , , , and filters along with BgVriz filters, we find that the ratio of peak pseudobolometric luminosity using BgVriz and SDAUBgVriz bands (, , , ) is , suggesting that the UV may be contributing up to % of the bolometric luminosity near peak. Using the same analytical model for the SDAUBgVriz pseudobolometric light curve, we find that M⊙ of 56Ni is synthesized during the explosion of SN 2024pxl. This is in agreement with our earlier estimate using optical data only. Based on the observed UV+optical data for SN 2024pxl, our final estimate for the 56Ni mass is M⊙.
Fink et al. (2014) presented numerical calculations for deflagration models of CO WDs with varying explosion energy (parametrized by the number of ignition kernels). These models span a broad range of explosion parameters (56Ni mass 0.03–0.38 M⊙, rise time 7.6–14.4 days). The estimated mass of 56Ni (0.03 M⊙) and rise time (10 days) for SN 2024pxl are consistent with the range provided by these models. However, the peak magnitude in the V band for SN 2024pxl ( mag) is on the fainter end of the estimation (16.84 to 18.96 mag) given by Fink et al. (2014). The mass of 56Ni in the case of SN 2024pxl matches with the 56Ni synthesized in the weakest-energy N1def (single ignition point) model presented by Fink et al. (2014). Radiative-transfer modeling based on the N1def model by Kwok et al. (2025) shows good agreement with the spectra of SN 2024pxl from the optical through the MIR. For low-luminosity Type Iax SNe, Kromer et al. (2015) studied the deflagration of a hybrid CONe WD. Our estimated peak magnitude of SN 2024pxl in the B band ( mag) and mass of 56Ni (0.03 M⊙) exceed the predictions from the hybrid CONe WD model ( to mag, mass of 56Ni M⊙) that may apply to lower-luminosity objects.
5 Spectral Evolution
5.1 Spectral Features and Comparisons
Figures 9, 10, and 11 present the spectral evolution (LABEL:tab:spectroscopic_observations_24pxl) of SN 2024pxl from approximately 8 to 89 days relative to Bmax. Spectra were obtained nearly daily for almost two months after maximum, producing a rich data set allowing detailed tracking of spectral evolution. We compare the spectral features of SN 2024pxl with those of other well-studied Type Iax SNe at similar pre-maximum, near-maximum, post-maximum, and late phases in Figures 12, 13, 14, and 15, respectively.
5.1.1 Pre-maximum spectra
We observe that at early, prepeak times, SN 2024pxl has a flat continuum consistent with the low temperatures we infer from the photometry (see Figure 8). The early-time spectra of SN 2024pxl display features such as Fe III, Fe II, and Si II (Figure 12). In agreement with our findings that SN 2024pxl is a photometrically intermediate Type Iax SN, it is spectroscopically intermediate as well, sharing similarities with both high-luminosity Type Iax SNe (e.g., SNe 2012Z, 2005hk, and 2020udy) and low-luminosity objects (e.g., SNe 2010ae and 2008ha). Specifically, SN 2024pxl has weak Si II features but also exhibits a C II line at 6580 Å. In high-luminosity Type Iax SNe, the Si II line is relatively weak, and C II is generally not detected. In contrast, in low-luminosity Type Iax SNe these features are much more prominent. The Ca II NIR triplet is present in SN 2024pxl but is weaker than in SNe 2010ae and 2008ha. This combination of weak Si II and presence of C II suggests that SN 2024pxl also serves as a spectroscopic transitional link between high-luminosity and low-luminosity Type Iax SNe.
5.1.2 Near-maximum spectra
The peak-brightness spectrum of SN 2024pxl (Figure 13) shows an increase in the strength of the Si II line at 6355 Å and the Fe III feature near 4400 Å. At this epoch, SN 2024pxl resembles SNe 2019muj (intermediate luminosity) and 2020rea (high luminosity). The Fe II features near 5000 Å are similar to those in SNe 2002cx, 2005hk (both high luminosity), and 2019muj, while features near 4000 Å match well with those of SN 2020rea. We detect the Ca II NIR triplet in SN 2024pxl, which is not detected in high-luminosity objects, but is even more pronounced in the low-luminosity SN 2010ae. The combination of Ca II NIR features similar to low/intermediate-luminosity objects, and Fe II characteristics shared with more luminous Type Iax SNe again implies that SN 2024pxl represents a transitional case.
5.1.3 Post-maximum and late-time spectra
Figure 14 shows the post-maximum spectra of SN 2024pxl in the context of other objects. At these epochs, SN 2024pxl shows remarkable similarity with SN 2019muj, an intermediate-luminosity SN Iax. During the post-maximum phase, the Si II and C II lines vanish, giving way to Fe and Co lines in the 5000-–7000 Å range. Iron-group elements (IGEs) and a strong Ca II NIR triplet dominate the post-maximum spectra, similar to other high- and intermediate-luminosity objects. Low-luminosity Type Iax SNe (e.g., SNe 2010ae and 2008ha, shown in Figure 14) evolve more rapidly and exhibit narrower spectral signatures than SN 2024pxl. The Cr II feature near 4800 Å and the Co II lines near 6500 Å in SN 2024pxl share strong resemblance with SNe 2002cx and 2019muj. Additionally, the Ca II NIR triplet in SN 2019muj is strikingly similar to that in SN 2024pxl.
We compare the late-time spectrum of SN 2024pxl obtained at 89 days with spectra of SNe 2008ge (Foley et al., 2010), 2014dt (Singh et al., 2018), and 2020udy at similar epochs (Figure 15). At late times, the spectral lines become narrower and several forbidden emission lines such as [Fe II], [Ni II], and [Ca II] emerge along with narrow permitted lines such as Fe II. The strength of these forbidden lines increases with the cooling of the ejecta. Owing to their intrinsically low luminosity, late-time spectra of Type Iax SNe are rare. However, the late-time spectra of Type Iax SNe vary widely, with some showing broad emission lines and others dominated by narrow features (Foley et al., 2016). Of the comparison SNe in Figure 15, SN 2024pxl most closely resembles SN 2014dt, exhibiting narrower lines than the higher-luminosity objects SNe 2020udy and 2008ge. The narrow features seen in the late-phase spectra of Type Iax SNe are thought to originate from the innermost regions (Maeda & Kawabata, 2022; Camacho-Neves et al., 2023). Additional late-time observations of Type Iax SNe across the luminosity distribution will help to understand the origin and nature of these spectral features more precisely.
Overall, at post-maximum and late times, SN 2024pxl shares similarities with high-luminosity Type Iax SNe and intermediate-luminosity Type Iax SN 2019muj.
5.1.4 NIR spectra
Figure 16 displays the NIR spectral evolution of SN 2024pxl from 6.7 to 42 days post maximum. The first two NIR spectra of SN 2024pxl show underdeveloped Fe II features with evidence for the Mg II 10,952 line. Signatures of C I 10,693 and C I 11,754 might be identifiable in the early NIR spectra of SN 2024pxl. As the ejecta cool and expand, the spectrum-formation region recedes to lower velocities and lines become somewhat less blended, with the Co II lines increasing in prominence over time. Our spectral sequence shows that Fe II and Co II lines dominate the NIR. These lines are distinctive features in the NIR spectra of all Type Iax SNe (Stritzinger et al., 2014), and Figure 17 displays this similarity in NIR spectra from bright to faint objects.
The NIR spectrum of SN 2024pxl at 23.4 days shows more similarities with the higher luminosity SN 2005hk (Kromer et al., 2013) than with the slightly less luminous SN 2019muj (Barna et al., 2021) and much less luminous SN 2010ae (Stritzinger et al., 2014). In particular, the Mg II absorption, the features at 12,000 Å and 12,700 Å, and the prominence of the Co II lines in the 16,000 Å region are differentiating features. Where SN 2024pxl resembles SN 2005hk, SN 2019muj instead resembles SN 2010ae, mainly owing to the lower velocities that allow individual lines to be separated more clearly. This NIR spectral comparison provides additional support that intermediate-luminosity objects of the class (like SNe 2024pxl and 2019muj) are transitional between high- and low-luminosity objects.
5.2 Velocity Evolution
We probe the evolution of velocities of C II, Si II, and Fe II lines using our optical spectral sequence, and the Co II lines from our NIR spectra. We measure line velocities by fitting Gaussian profiles to the absorption trough of the P-Cygni profiles of the C II 6580, Si II 6355, Fe II 5156, Co II 15,759, Co II 16,064, and Co II 16,361 lines (Figure 18). This standard method of measuring photospheric line velocities results in Co II measurements that are consistent with those presented by Kwok et al. (2025), obtained by fitting blended Co II lines simultaneously. Figure 19 shows the Si II line-velocity evolution for SN 2024pxl alongside several other well-studied Type Iax SNe. The uncertainties presented in Figure 19 could likely be underestimated, as we only account for measurement uncertainties from our fitting (and not for the effects of line blending, for example).
SN 2024pxl exhibits lower velocities than other SNe shown in Figure 19 except SN 2008ha. The prominent Si II line is overlapped by a continuously growing Fe II feature in the red wing of the Si II line. As a result, we systematically underestimate the velocities associated with this line over time by considering it as a single absorption feature. In the case of SN 2020udy, Maguire et al. (2023) showed that the Si II line is contaminated by the Fe II line before maximum light, although Si II is present in post-maximum spectra. This might be why a significant drop is observed in the Si II velocities post-maximum in SN 2024pxl. It is likely that more careful modeling of these lines will be required to obtain a more accurate comparison of the Si II line-velocity evolution across Type Iax SNe.
The Si II and C II line velocities of SN 2024pxl at peak are 3800 km s-1 and 3500 km s-1, respectively. The measured Fe II velocities are 7100, 7060, and 5950 km s-1 at epochs , 5.8, and 1.3 days since Bmax, respectively. At a similar epoch, the estimated velocities using Fe II lines are higher than the Si II and C II line velocities, indicating significant mixing in the burning products (Phillips et al., 2007). Kwok et al. (2025) also find evidence of mixing in SN 2024pxl from the emission lines of IGEs and intermediate-mass elements (IMEs), which are all centrally peaked and have similar velocity offsets and widths. The measured Co II line velocities in the NIR region for SN 2024pxl at an epoch of 23.4 days are 3030, 2730, and 2750 km s-1 for Co II 15,759, Co II 16,064, and Co II 16,361, respectively (consistent with one another), and slightly higher than those reported by Barna et al. (2021) at an epoch of 26.5 days past Bmax for SN 2019muj. At a comparable epoch, velocities of Co II lines in SN 2010ae (Stritzinger et al., 2014) are lower than in SN 2024pxl.
6 Summary
SN 2024pxl is a nearby Type Iax SN with an early detection that enabled precise constraints on its rise time. With a peak absolute magnitude of mag, SN 2024pxl is an intermediate-luminosity Type Iax SN similar to SN 2019muj. The color evolution of SN 2024pxl aligns closely with that of other Type Iax SNe. Utilizing Swift UV coverage around peak brightness, we estimate the contribution of UV flux to the bolometric luminosity and find that SN 2024pxl’s bolometric luminosity falls between those of SNe 2002cx and 2019muj, placing it in a unique intermediate position in the luminosity range of Type Iax SNe. Analytical modeling of the pseudobolometric light curve indicates the synthesis of M⊙ of 56Ni and an ejected mass of M⊙. The explosion parameters calculated through analytical modeling of the pseudobolometric light curve are consistent with the N1def model from Fink et al. (2014), a weak deflagration of a near- WD with one ignition point. This is consistent with the results of Kwok et al. (2025) that SN 2024pxl displays spectroscopic similarity to the predictions of the N1def model.
Our spectral comparisons also suggest that SN 2024pxl is a transitional object between high-luminosity and low-luminosity Type Iax SNe. It exhibits many similarities to SN 2019muj and high-luminosity Type Iax SNe. In the early spectral sequence, we also detect C II that is seen more prominently in less luminous objects. Additionally, the late-time spectra of SN 2024pxl show narrow emission lines around 7300 Å whereas the high-luminosity objects (e.g., SNe 2008ge and 2020udy) display broader features.
In the NIR, SN 2024pxl shows signatures from Mg II, Fe II, and Co II and it resembles the high-luminosity Type Iax SN 2005hk, while SN 2019muj more closely resembles the low-luminosity Type Iax SN 2010ae. This is mainly due to more line blending in SN 2024pxl compared with SN 2019muj, caused by wider line widths. Despite the higher velocities indicated by the wider line widths in SN 2024pxl, the evolution of the photospheric velocity of SN 2024pxl shows systematically lower velocities relative to other comparison SNe, except for the very low-luminosity SN 2008ha. Kwok et al. (2025) find potential evidence for bulk motion on the order of a few 100 km s-1 in the ejecta of SN 2024pxl from measurements of MIR forbidden lines, which would impact the true vs. observed photospheric velocity. Given the spectroscopically and photometrically intermediate position of SN 2024pxl between SN 2019muj and brighter Type Iax SNe, we suggest that velocity measurements of the widths of the Co II lines may be better correlated with luminosity than the photospheric velocity. These spectral properties provide evidence that SN 2024pxl is a transitional Type Iax SN.
We measure the Fe II velocity to be substantially higher than that of Si II and C II at similar epochs. At a slightly later phase, Co II velocity measurements seem to align with the expected velocity decay from the Fe II measurements. Higher velocities of IGEs compared to IMEs indicate mixing in the ejecta, consistent with the findings of Kwok et al. (2025).
Overall, SN 2024pxl is a compelling link between high-luminosity and low-luminosity Type Iax SNe. The analysis of UV, optical, and NIR data presented in this work, and the spectral analysis of JWST NIR and MIR data by Kwok et al. (2025), demonstrate that SN 2024pxl is photometrically and spectroscopically consistent with the one-ignition-point weak deflagration model, N1def, from Fink et al. (2014). Kwok et al. (2025) also show that SN 2024pxl shares spectroscopic similarities (aside from the large luminosity difference) with the very faint Type Iax SN 2024vjm. These findings indicate that intermediate-luminosity objects such as SNe 2019muj and 2024pxl are links in a continuous distribution. Future studies of other Type Iax SNe with rich datasets, such as that presented for SN 2024pxl, from both ground- and space-based telescopes will help us better understand the nature of these peculiar SNe.
Remarkably, the diverse class of Type Iax SNe can span a huge luminosity range. Still, the seemingly continuous properties we observe may yet point to a common (or at least quite similar) progenitor origin and explosion mechanism. This contrasts with normal Type Ia SNe, for which multiple progenitor pathways may be plausible while producing more homogeneous outcomes. Further study of peculiar thermonuclear SNe holds promise to enable a better understanding of which WDs may explode and their explosion mechanisms.
7 acknowledgments
We thank Stéphane Blondin for insightful and helpful comments which improved the manuscript.
This work is based in part on observations collected at the European Southern Observatory under ESO programme 114.27JL.001. Partly based on observations made with the Nordic Optical Telescope, owned in collaboration by the University of Turku and Aarhus University. This work also makes use of data gathered with the 6.5 m Magellan telescopes at Las Campanas Observatory, Chile. A FIRE spectrum was obtained through A.P.’s prior support by a Carnegie Fellowship. This work makes use of data from the Las Cumbres Observatory global network of telescopes. The LCO group is supported by National Science Foundation (NSF) grants AST-1911151 and AST-1911225.
The SALT observations were obtained with Rutgers University program 2024-1-MLT-004 (PI L. A. Kwok). We are grateful to SALT Astronomer Rosalind Skelton for taking these data. Data reported here were obtained in part at the MMT Observatory, a joint facility of the University of Arizona and the Smithsonian Institution. We sincerely thank the MMT observers and staff for their accommodation of our strict timing requirements to ensure they coincided with JWST observations.
Some of the data presented herein were obtained at the W. M. Keck Observatory, which is operated as a scientific partnership among the California Institute of Technology, the University of California, and the National Aeronautics and Space Administration (NASA). The Observatory was made possible by the generous financial support of the W. M. Keck Foundation. This work was supported by a NASA Keck PI Data Award (PI D. J. Sand), administered by the NASA Exoplanet Science Institute. The authors wish to recognize and acknowledge the very significant cultural role and reverence that the summit of Maunakea has always had within the Indigenous Hawaiian community. We are most fortunate to have the opportunity to conduct observations from this mountain.
A major upgrade of the Kast spectrograph on the Shane 3 m telescope at Lick Observatory, led by Brad Holden, was made possible through generous gifts from the Heising-Simons Foundation, William and Marina Kast, and the University of California Observatories. Research at Lick Observatory is partially supported by a generous gift from Google. We would like to express our gratitude to the Lick Observatory staff for their support. Shane 3 m observations were conducted on the stolen land of the Ohlone (Costanoans), Tamyen and Muwekma Ohlone tribes. We gratefully acknowledge the usage of native lands for our science. Observations from coauthor A.J.M. were made under the aegis of the ASTRAL (Astronomy/STEM Alliance with Lick Observatory) consortium, supported by a generous grant from the Gordon and Betty Moore Foundation (PI B. Macintosh).
We thank the staff of IAO, Hanle, CREST, and Hosakote, who made these observations possible. The facilities at IAO and CREST are operated by the Indian Institute of Astrophysics, Bangalore. We thank the Subaru staff for the data taken by the Subaru Telescope (S23A-023). Based in part on observations made with the Gran Telescopio Canarias (GTC), installed at the Spanish Observatorio del Roque de los Muchachos of the Instituto de Astrofísica de Canarias, on the island of La Palma, using OSIRIS and EMIR instruments.
Based (in part) on data acquired at the ANU 2.3 m telescope. The automation of the telescope was made possible through an initial grant provided by the Centre of Gravitational Astrophysics and the Research School of Astronomy and Astrophysics at the Australian National University and through a grant provided by the Australian Research Council through LE230100063. We acknowledge the traditional custodians of the land on which the telescope stands, the Gamilaraay people, and pay our respects to elders past and present.
The HET is a joint project of the University of Texas at Austin, the Pennsylvania State University, Stanford University, Ludwig-Maximilians-Universität München, and Georg-August-Universität Göttingen. The HET is named in honor of its principal benefactors, William P. Hobby and Robert E. Eberly. The Marcario Low-Resolution Spectrograph is named for Mike Marcario of High Lonesome Optics who fabricated several optics for the instrument but died before its completion. The LRS is a joint project of the HET partnership and the Instituto de Astronomía de la Universidad Nacional Autónoma de México.
We acknowledge Weizmann Interactive Supernova data REPository http://wiserep.weizmann.ac.il (WISeREP) (Yaron & Gal-Yam, 2012). This research has made use of the CfA Supernova Archive, which is funded in part by the NSF through grant AST-0907903. This research has made use of the NASA/IPAC Extragalactic Database (NED) which is operated by the Jet Propulsion Laboratory, California Institute of Technology, under contract with NASA.
M.S. acknowledges financial support provided under the National Post Doctoral Fellowship (N-PDF; File Number: PDF/2023/002244) by the Science & Engineering Research Board (SERB), Anusandhan National Research Foundation (ANRF), Government of India. L.A.K. is supported by a CIERA Postdoctoral Fellowship.
Support for this research at Rutgers University (S.W.J., C.L., M.S.) was provided by NSF award AST-2407567.
A.A.M., C.L., and N.R. are supported by DoE award #DE-SC0025599. MMT and Keck Observatory access for N.R. and C.L. was supported by Northwestern University and the Center for Interdisciplinary Exploration and Research in Astrophysics (CIERA). A.C.G. and the Fong Group at Northwestern acknowledge support by the NSF under grants AST-1909358, AST-2206494, AST-2308182, and CAREER grant AST-2047919.
G.C.A. thanks the Indian National Science Academy for support under the INSA Senior Scientist Programme. A.F. acknowledges support by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (ERC Advanced Grant KILONOVA #885281) and the State of Hesse within the Cluster Project ELEMENTS. C.L. acknowledges support from DOE award DE-SC0010008 to Rutgers University. M.R.S. is supported by the STScI Postdoctoral Fellowship.
Time-domain research by the University of Arizona team and D.J.S. is supported by NSF grants 2108032, 2308181, 2407566, and 2432036 and the Heising-Simons Foundation under grant #2020-1864. Time-domain research by the University of California, Davis team and S.V. is supported by NSF grant AST-2407565. K.A.B. is supported by an LSST-DA Catalyst Fellowship; this publication was thus made possible through the support of grant 62192 from the John Templeton Foundation to LSST-DA. N.F. acknowledges support from the NSF Graduate Research Fellowship Program under grant DGE-2137419.
J.E.A. is supported by the international Gemini Observatory, a program of NSF’s NOIRLab, which is managed by the Association of Universities for Research in Astronomy (AURA) under a cooperative agreement with the NSF, on behalf of the Gemini partnership of Argentina, Brazil, Canada, Chile, the Republic of Korea, and the United States of America. J.A.V. acknowledges the Postgraduate School of the Universidad de Antofagasta for its support and allocated grants. R.D. acknowledges funds by ANID grant FONDECYT Postdoctorado #3220449.
A.V.F.’s group at UC Berkeley received financial assistance from the Christopher R. Redlich Fund, as well as donations from Gary and Cynthia Bengier, Clark and Sharon Winslow, Alan Eustace and Kathy Kwan, William Draper, Timothy and Melissa Draper, Briggs and Kathleen Wood, Sanford Robertson (W.Z. is a Bengier-Winslow-Eustace Specialist in Astronomy, T.G.B. is a Draper-Wood-Robertson Specialist in Astronomy, Y.Y. was a Bengier-Winslow-Robertson Fellow in Astronomy), and numerous other donors.
K. Maguire acknowledges funding from Horizon Europe ERC grant 101125877. J.H.T. acknowledges support from EU H2020 ERC grant 758638. T.T. acknowledges support from NSF grant AST-2205314 and the NASA ADAP award 80NSSC23K1130. K. Maede acknowledges support from JSPS KAKENHI grants JP24KK0070, JP24H01810, and JP20H00174, and from JSPS Bilateral Joint Research Project (JPJSBP120229923). K. Misra acknowledges support from the BRICS grant DST/ICD/BRICS/Call-5/CoNMuTraMO/2023 (G) funded by the DST, India. J.V. is supported by NKFIH-OTKA grant K142534. G.C.A. thanks the Indian National Science Academy for support under the INSA Senior Scientist Programme. M.R.S. is supported by the STScI Postdoctoral Fellowship. B.B. received support from the Hungarian National Research, Development and Innovation Office grants OTKA PD-147091. L.G. acknowledges financial support from AGAUR, CSIC, MCIN and AEI 10.13039/501100011033 under projects PID2023-151307NB-I00, PIE 20215AT016, CEX2020-001058-M, ILINK23001, COOPB2304, and 2021-SGR-01270. H.K. was funded by the Research Council of Finland projects 324504, 328898, and 353019. D.A.H., G.H., and C.M. were supported by NSF grants AST-1313484 and AST-1911225.
Appendix A Optical photometry and spectroscopy of SN 2017drh
We present the optical Las Cumbres Observatory photometry and FLOYDS spectroscopy for SN 2017drh, which exploded in the same host galaxy, NGC 6384, as SN 2024pxl. The spectra were reduced using the FLOYDS reduction pipeline (Valenti et al., 2014). The photometric observations were taken as part of the Global Supernova Project (GSP) collaboration. These data were reduced with lcogtsnpipe (Valenti et al., 2016b), a PyRAF-based image-reduction pipeline. In Figure 20, the LCO data are shown with the BayeSN light curve fits that also included data from Stahl et al. (2019).
| Date | JD† | Phase‡ | ||||||
| (Days) | (mag) | (mag) | (mag) | (mag) | (mag) | (mag) | ||
| 2017-05-10 | 883.54 | 7.66 | 18.510.26 | 17.460.03 | — | — | — | — |
| 2017-05-10 | 883.55 | 7.65 | — | 17.520.05 | 17.160.02 | 16.350.02 | — | — |
| 2017-05-10 | 883.56 | 7.64 | — | — | 17.150.02 | — | 15.740.01 | 15.270.02 |
| 2017-05-13 | 887.48 | 3.72 | 18.370.08 | — | — | — | — | — |
| 2017-05-13 | 887.49 | 3.71 | 18.410.08 | 17.090.03 | — | — | — | — |
| 2017-05-14 | 887.50 | 3.70 | — | 17.070.04 | 16.730.02 | 15.870.02 | — | — |
| 2017-05-14 | 887.51 | 3.69 | — | — | — | — | 15.350.01 | 14.990.01 |
| 2017-05-15 | 888.53 | 2.67 | 18.010.21 | — | — | — | — | — |
| 2017-05-15 | 888.54 | 2.66 | 17.960.19 | 16.990.04 | — | — | — | — |
| 2017-05-15 | 888.55 | 2.65 | — | — | 16.590.04 | 15.980.02 | — | — |
| 2017-05-15 | 888.56 | 2.64 | — | — | — | — | 15.250.02 | 14.970.03 |
| 2017-05-16 | 889.56 | 1.64 | 18.690.09 | — | — | — | — | — |
| 2017-05-16 | 890.47 | 0.73 | 18.320.07 | — | — | — | — | — |
| 2017-05-16 | 890.48 | 0.72 | — | 16.950.03 | — | 15.750.02 | — | — |
| 2017-05-18 | 891.50 | 0.30 | 18.410.06 | — | — | — | — | — |
| 2017-05-18 | 891.51 | 0.31 | 18.400.06 | 17.060.02 | — | 15.860.02 | — | — |
| 2017-05-18 | 891.52 | 0.32 | — | — | 16.630.01 | 15.810.01 | 15.280.01 | — |
| 2017-05-18 | 891.53 | 0.33 | — | — | — | — | 15.280.01 | 15.160.01 |
| 2017-05-18 | 892.45 | 1.25 | 18.110.19 | — | — | — | — | — |
| 2017-05-18 | 892.46 | 1.26 | — | 16.980.02 | — | 15.760.01 | — | — |
| 2017-05-18 | 892.47 | 1.27 | — | — | 16.610.01 | — | 15.260.01 | — |
| 2017-05-18 | 892.48 | 1.28 | — | — | — | — | — | 15.150.01 |
| 2017-05-21 | 895.49 | 4.29 | 18.210.18 | — | — | — | — | — |
| 2017-05-22 | 895.50 | 4.30 | 18.230.18 | 17.210.03 | — | — | — | — |
| 2017-05-22 | 895.51 | 4.31 | — | — | 16.690.01 | 15.800.02 | — | — |
| 2017-05-22 | 895.52 | 4.32 | — | — | — | — | 15.310.01 | 15.240.01 |
| 2017-05-22 | 896.19 | 4.99 | 18.760.07 | — | — | — | — | — |
| 2017-05-22 | 896.20 | 5.00 | 18.700.08 | 17.080.02 | — | — | — | — |
| 2017-05-22 | 896.21 | 5.01 | — | 17.230.02 | 16.630.01 | 15.840.02 | — | — |
| 2017-05-22 | 896.22 | 5.02 | — | — | 16.620.01 | — | 15.270.01 | 15.230.01 |
| 2017-05-23 | 897.19 | 5.99 | 18.550.21 | — | — | — | — | — |
| 2017-05-23 | 897.20 | 6.00 | — | 17.450.03 | — | 16.020.02 | — | — |
| 2017-05-23 | 897.21 | 6.01 | — | — | 16.710.01 | 15.860.01 | 15.320.05 | — |
| 2017-05-23 | 897.22 | 6.02 | — | — | — | — | 15.330.01 | — |
| 2017-05-26 | 900.20 | 9.00 | 19.210.08 | — | — | — | — | — |
| 2017-05-26 | 900.21 | 9.01 | 18.980.32 | 17.450.03 | — | — | — | — |
| 2017-05-26 | 900.22 | 9.02 | — | — | 16.820.01 | 16.070.02 | — | — |
| 2017-05-26 | 900.23 | 9.03 | — | — | — | — | 15.550.01 | 15.530.01 |
| 2017-05-31 | 905.18 | 13.98 | 19.080.34 | — | — | — | — | — |
| 2017-05-31 | 905.19 | 13.99 | — | 18.380.04 | — | — | — | — |
| 2017-05-31 | 905.20 | 14.00 | — | — | 17.330.02 | 16.420.02 | — | — |
| 2017-05-31 | 905.21 | 14.01 | — | — | — | — | 15.810.01 | 15.540.01 |
| 2017-06-05 | 910.37 | 19.17 | 20.210.35 | — | — | — | — | — |
| 2017-06-05 | 910.38 | 19.18 | 20.370.30 | 18.640.12 | — | — | — | — |
| 2017-06-05 | 910.39 | 19.19 | — | — | 18.000.04 | 16.610.03 | — | — |
| 2017-06-05 | 910.40 | 19.20 | — | — | — | — | 15.960.02 | 15.510.02 |
| 2017-06-09 | 913.73 | 22.53 | — | — | — | — | 16.040.03 | — |
| 2017-06-09 | 913.74 | 22.54 | — | — | — | 16.910.04 | — | — |
| 2017-06-09 | 913.75 | 22.55 | — | 18.970.22 | — | 16.800.04 | — | — |
| 2017-06-09 | 913.76 | 22.56 | — | 19.210.25 | — | — | — | — |
| 2017-06-15 | 919.88 | 28.68 | — | — | 18.720.07 | — | 16.530.03 | — |
| 2017-06-15 | 919.89 | 28.69 | — | 18.760.23 | 18.960.15 | — | — | — |
| 2017-06-15 | 920.06 | 28.86 | — | — | 18.830.06 | 17.490.04 | — | — |
| 2017-06-15 | 920.07 | 28.87 | — | — | — | — | 16.430.04 | — |
| 2017-06-15 | 920.08 | 28.88 | — | — | 18.750.09 | — | — | — |
| 2017-06-15 | 920.09 | 28.89 | — | 19.560.18 | — | 17.510.04 | — | — |
| 2017-06-15 | 920.10 | 28.90 | 20.630.35 | — | — | — | — | — |
| 2017-06-15 | 920.11 | 28.91 | — | — | — | — | 16.550.02 | 15.870.02 |
| 2017-06-22 | 927.07 | 35.87 | — | 19.710.07 | — | — | — | — |
| 2017-06-22 | 927.08 | 35.88 | — | 19.820.07 | — | 17.910.05 | — | 16.310.03 |
| 2017-06-22 | 927.09 | 35.89 | — | — | 19.040.06 | — | 16.940.06 | — |
| 2017-06-26 | 931.11 | 39.91 | 20.980.48 | — | — | — | — | — |
| 2017-06-26 | 931.12 | 39.92 | 20.250.43 | — | — | — | — | — |
| 2017-06-26 | 931.13 | 39.93 | — | 19.640.23 | 19.070.11 | 17.830.09 | — | — |
| 2017-06-26 | 931.14 | 39.94 | — | — | 18.980.07 | — | 17.050.03 | 16.480.06 |
| 2017-06-26 | 931.15 | 39.95 | — | — | — | — | — | 16.610.12 |
| 2017-06-28 | 932.83 | 41.63 | 19.720.43 | — | — | — | — | — |
| 2017-06-28 | 932.84 | 41.64 | 19.930.54 | 19.690.18 | — | — | — | — |
| 2017-06-28 | 932.85 | 41.65 | — | 19.800.17 | — | 17.840.05 | — | 16.620.03 |
| 2017-06-28 | 932.86 | 41.66 | — | — | — | — | — | 16.630.02 |
| 2017-06-30 | 934.85 | 43.65 | — | — | — | 18.000.03 | — | — |
| 2017-07-01 | 935.61 | 44.41 | 21.060.26 | — | — | — | — | — |
| 2017-07-01 | 935.62 | 44.42 | 21.010.21 | 19.940.13 | — | — | — | — |
| 2017-07-01 | 935.63 | 44.43 | — | 19.860.16 | 19.380.07 | 18.000.05 | — | — |
| 2017-07-01 | 935.64 | 44.44 | — | — | 19.410.07 | — | 17.400.03 | 16.850.03 |
| 2017-07-01 | 935.65 | 44.45 | — | — | — | — | — | 16.850.03 |
| 2017-07-03 | 938.43 | 47.23 | — | 19.850.19 | — | — | — | — |
| 2017-07-03 | 938.44 | 47.24 | — | 20.250.17 | — | 17.980.05 | — | — |
| 2017-07-03 | 938.45 | 47.25 | — | — | 19.500.10 | — | 17.490.03 | — |
| 2017-07-03 | 938.46 | 47.26 | — | — | — | — | 17.470.03 | 16.900.03 |
| 2017-07-09 | 943.67 | 52.47 | — | 20.390.35 | — | — | — | — |
| 2017-07-09 | 943.68 | 52.48 | — | — | 19.120.26 | 18.230.08 | — | — |
| 2017-07-09 | 943.69 | 52.49 | — | — | 18.650.22 | — | 17.420.14 | — |
| 2017-07-09 | 943.70 | 52.50 | — | — | — | — | — | 16.930.12 |
| 2017-07-15 | 950.05 | 58.85 | — | 19.800.19 | — | — | — | — |
| 2017-07-15 | 950.06 | 58.86 | — | — | 19.590.12 | 18.240.06 | — | — |
| 2017-07-15 | 950.07 | 58.87 | — | — | 19.510.11 | — | 17.780.03 | — |
| 2017-07-15 | 950.08 | 58.88 | — | — | — | — | — | 17.320.04 |
| 2017-07-21 | 956.29 | 65.09 | — | 20.260.08 | — | — | — | — |
| 2017-07-21 | 956.30 | 65.10 | — | 20.290.08 | — | 18.500.03 | — | — |
| 2017-07-21 | 956.31 | 65.11 | — | — | 19.780.11 | — | 18.120.05 | — |
| 2017-07-21 | 956.32 | 65.12 | — | — | — | — | 18.140.05 | 17.590.05 |
| 2017-07-27 | 962.34 | 71.14 | — | 20.390.10 | — | — | — | — |
| 2017-07-27 | 962.35 | 71.15 | — | — | 19.740.09 | 18.650.04 | — | — |
| 2017-07-27 | 962.36 | 71.16 | — | — | 19.700.10 | — | 18.310.05 | — |
| 2017-07-27 | 962.37 | 71.17 | — | — | — | — | — | 17.820.06 |
| 2017-08-02 | 968.32 | 77.12 | — | 20.650.38 | — | — | — | — |
| 2017-08-02 | 968.33 | 77.13 | — | 20.880.48 | — | 18.650.11 | — | — |
| 2017-08-02 | 968.34 | 77.14 | — | — | 20.090.18 | — | 18.470.09 | — |
| 2017-08-02 | 968.35 | 77.15 | — | — | — | — | 18.590.10 | — |
| 2017-08-08 | 973.98 | 82.78 | — | 19.860.21 | — | — | — | — |
| 2017-08-08 | 973.99 | 82.79 | — | — | 19.530.13 | 19.360.18 | — | — |
| 2017-08-08 | 974.00 | 82.80 | — | — | 20.040.14 | — | 18.500.11 | — |
| 2017-08-08 | 974.01 | 82.81 | — | — | — | — | — | 18.160.11 |
| 2017-08-14 | 979.98 | 88.78 | — | 20.150.21 | — | — | — | — |
| 2017-08-14 | 979.99 | 88.79 | — | 20.580.14 | — | 18.960.10 | — | — |
| 2017-08-14 | 980.00 | 88.80 | — | — | 19.180.26 | — | — | — |
| 2017-08-19 | 984.55 | 93.35 | — | 20.030.14 | — | — | — | — |
| 2017-08-19 | 984.56 | 93.36 | — | 20.660.18 | — | 19.270.10 | — | — |
| 2017-08-19 | 984.57 | 93.37 | — | — | 19.980.09 | — | 19.240.10 | — |
| 2017-08-19 | 984.58 | 93.38 | — | — | — | — | 19.310.11 | 18.600.10 |
| 2017-08-25 | 990.56 | 99.36 | — | 20.260.18 | — | 19.380.08 | — | — |
| 2017-08-25 | 990.57 | 99.37 | — | — | 20.020.11 | 19.520.07 | — | — |
| 2017-08-25 | 990.58 | 99.38 | — | — | — | — | 19.260.10 | 18.790.10 |
| 2017-08-25 | 990.59 | 99.39 | — | — | — | — | — | 18.880.09 |
| 2017-08-31 | 996.94 | 105.74 | — | 19.530.09 | — | 19.320.09 | — | — |
| 2017-08-31 | 996.95 | 105.75 | — | — | 19.860.10 | 20.440.50 | — | — |
| 2017-08-31 | 996.96 | 105.76 | — | — | — | — | 19.350.09 | 19.090.11 |
| 2017-08-31 | 996.97 | 105.77 | — | — | — | — | — | 19.930.22 |
| 2017-09-06 | 1002.92 | 111.72 | — | 21.220.32 | — | — | — | — |
| 2017-09-06 | 1002.93 | 111.73 | — | — | 20.370.13 | 19.720.25 | — | — |
| 2017-09-06 | 1002.94 | 111.74 | — | — | 20.610.19 | — | 20.000.16 | — |
| 2017-09-06 | 1002.95 | 111.75 | — | — | — | — | — | 19.100.16 |
| 2017-09-12 | 1008.63 | 117.43 | — | 20.830.15 | — | — | — | — |
| 2017-09-12 | 1008.64 | 117.44 | — | — | 20.070.13 | 19.870.18 | — | — |
| 2017-09-12 | 1008.65 | 117.45 | — | — | 20.500.17 | — | 19.870.13 | — |
| 2017-09-12 | 1008.66 | 117.46 | — | — | — | — | — | 19.240.13 |
| 2017-09-16 | 1012.88 | 121.68 | — | 20.350.21 | — | — | — | — |
| 2017-09-16 | 1012.89 | 121.69 | — | — | 21.080.14 | 19.820.14 | — | — |
| 2017-09-16 | 1012.90 | 121.70 | — | — | 20.070.11 | — | — | — |
| 2017-09-16 | 1012.91 | 121.71 | — | — | — | — | 20.180.20 | 19.530.17 |
| 2017-09-16 | 1012.92 | 121.72 | — | — | — | — | — | 19.460.17 |
| 2017-10-09 | 1035.59 | 144.39 | — | 21.340.17 | — | — | — | — |
| 2017-10-09 | 1035.60 | 144.40 | — | — | — | 20.290.10 | — | — |
| 2017-10-09 | 1035.61 | 144.41 | — | — | 20.780.12 | — | — | — |
| 2017-10-09 | 1035.62 | 144.42 | — | — | — | — | 20.550.19 | 20.100.25 |
| 2017-10-09 | 1035.63 | 144.43 | — | — | — | — | — | 21.800.48 |
| 2017-10-17 | 1043.56 | 152.36 | — | 21.720.40 | — | — | — | — |
| 2017-10-17 | 1043.57 | 152.37 | — | 21.470.25 | — | 20.610.15 | — | — |
| 2017-10-17 | 1043.58 | 152.38 | — | — | — | 20.190.12 | — | — |
| 2017-10-17 | 1043.59 | 152.39 | — | — | 21.670.14 | — | 20.370.21 | — |
| 2017-10-17 | 1043.60 | 152.40 | — | — | — | — | 20.610.19 | 19.870.22 |
| 2018-03-07 | 1184.92 | 293.72 | — | 21.150.44 | — | — | — | — |
| 2018-03-07 | 1184.93 | 293.73 | — | 20.960.42 | — | — | — | — |
| 2018-03-07 | 1184.94 | 293.74 | — | — | — | 21.460.45 | — | — |
| 2018-03-07 | 1184.95 | 293.75 | — | — | 22.270.30 | 21.190.41 | — | — |
| 2018-03-07 | 1184.96 | 293.76 | — | — | 23.120.31 | — | — | — |
| 2018-03-07 | 1184.97 | 293.77 | — | — | — | — | 22.140.16 | — |
| 2018-03-07 | 1184.98 | 293.78 | — | — | — | — | — | 21.410.23 |
| 2018-03-07 | 1184.99 | 293.79 | — | — | — | — | — | 21.700.24 |
| † JD 2,457,000+ . ‡Phase calculated with respect to . | ||||||||
| Date | JD† | Phase‡ | Telescope/Instrument |
| 2017-05-03 | 877.07 | 14.12 | FTN/FLOYDS |
| 2017-05-09 | 883.19 | 8.00 | FTN/FLOYDS |
| 2017-05-15 | 889.21 | 1.98 | FTN/FLOYDS |
| 2017-05-19 | 893.05 | 1.86 | FTN/FLOYDS |
| 2017-05-25 | 899.06 | 7.86 | FTN/FLOYDS |
| 2017-05-31 | 905.04 | 13.85 | FTN/FLOYDS |
| 2017-06-06 | 911.06 | 19.86 | FTN/FLOYDS |
| 2017-06-27 | 932.01 | 40.81 | FTN/FLOYDS |
†JD 2,457,000+ . ‡ Phase calculated with respect to .